# Continual Learning of Large Language Models: A Comprehensive Survey

### v1

@@ 4.1.1 CPT: Effectiveness and Efficiency
4.1.2 General Observations on CPT
The analysis presented in Table 1 sheds light on the prevailing research trends in continual pre-training (CPT). Firstly, it is evident that the development of advanced techniques tailored specifically for CPT is still at the starting stage and warrants further exploration. This observation is underscored by the fact that only about half of the examined papers propose novel techniques (9 out of 16 papers, represented in the deep gray section of Table 1), while the remaining half either focus solely on the effects of pure adaptation without considering continual learning techniques (3 out of 16 papers, represented in the white section), or conduct empirical studies on the straightforward application of existing continual learning techniques (4 out of 16 papers, represented in the light gray section). Secondly, while research extensively covers various continual learning techniques, such as rehearsal, parameter regularization, and architecture expansion (as indicated in the light gray section of Table 1), the practical incorporation of these techniques in systems remains relatively limited. Most practical implementations primarily focus on architecture expansion of LLMs [6, 55, 91, 68, 231, 46], with only a few explicitly utilizing replay [231, 46] and parameter regularization [6, 46] explicitly (deep gray section of Table 1). Thirdly, there is a pressing need for exploration into longer sequences of incremental phases in continual pre-training. Currently, the longest sequence of domains explored is eight, with content-level distributional shifts [127, 91]. However, this falls short of real-world scenarios where continual pre-training may occur more frequently and persist for months or years. The efficacy of continual learning techniques in such prolonged scenarios remains uncertain, as potential performance degradation with longer domain sequences is observed in techniques like EWC [140]. Additionally, investigating CPT in a task-boundary-free data stream setting is an important avenue for research as well.
4.1.3 Distributional Shifts in CPT
This survey categorizes distributional shifts of continual pre-training into three main types: (i) Language Shift: LLMs sequentially learn different language corpora, e.g., English $\rightarrow$ Chinese, focusing on token and vocabulary distributional shifts [82, 149]. (ii) Content Shift: LLMs sequentially learn corpora from different fields, e.g., chemistry $\rightarrow$ biology, focusing on token and vocabulary distributional shifts as well as shift of semantic meaning [91, 55, 127, 231, 46, 89]. (iii) Temporal Shift: Distributional shifts occur over time, e.g., news in 2021 $\rightarrow$ news in 2022, focusing on token and vocabulary shifts, and timestamp-sensitive knowledge retention and update, which aligns with real-world LLM deployment needs [6, 127, 68, 120, 119].
Language Shift. In contrast to the common approach of pre-training multilingual language models jointly on large corpora from multiple languages, [82] focuses on assessing these models’ natural ability to learn new languages sequentially (English, Norwegian, and Icelandic). The study does not employ explicit continual learning techniques for preventing horizontal forgetting. Nevertheless, it observes consistent positive forward transfer, facilitating new language acquisition regardless of the presentation order. Forgetting, on the other hand, emerges as a significant challenge, influenced by language order and not mitigated by increasing LLM size. In [149], the degree of forgetting of previously learned language (English) when adapting LLMs to a new language (Traditional Chinese) is investigated. Various continual learning techniques, including parameter freezing, LoRA [109], and (IA)3 [168], are evaluated across multiple dimensions, including output language, general knowledge retention, and reliability. Preliminary experimental results presented in this study highlight the non-trivial nature of addressing horizontal forgetting in continually pre-training LLMs under the language shift. To summarize, research on continual pre-training for language shifts is in its preliminary stages for two main reasons: Firstly, the datasets’ scale, including the number of languages and total token count, remains small. Secondly, specific methods targeting language shifts have yet to be proposed; only basic combinations of existing continual learning techniques have been evaluated.
Content Shift. Without using complex CL techniques, [89] continues the pre-training phase of Pythia [21] on the newly collected SlimPajama dataset [266]. The study focuses on optimizing continual pre-training by learning rate (re)warm-up. They discover that regardless of whether a larger or smaller maximum learning rate is used, models that undergo re-warming consistently exhibit improvements over models trained from scratch, even in terms of adaptation solely.
Another pioneering work, LLPT [127], establishes a comprehensive training and evaluation protocol for a series of content-level distributional shifts, referred to as “domain-incremental data streams” in the paper. They assess multiple continual learning methods based on masked language modeling perplexity for pre-training tasks and downstream task accuracy. Similar to findings in [82], they note that later domains benefit from knowledge learned from earlier ones, yet horizontal forgetting remains a significant challenge for earlier domains. Contrary to the common belief that experience replay (ER, [38]) is the most efficient approach to preventing forgetting, the authors find it scarcely improves continual pre-training performance. They speculate that ER’s inefficiency may stem from overfitting issues, as replaying with distillation loss can alleviate this problem efficiently [340, 127]. Following LLPT, Recyclable Tuning [231] is the first study to consider both upstream LLM suppliers and downstream consumers at the same time. It shows that if the upstream supplier continually pre-trains LLMs – initializing from the previous checkpoint and continuing pre-training on newly collected data, with or without replay, consumer-side efficiency can be boosted by recycling previously learned incremental components. Two CL techniques, initializing from outdated components and knowledge distillation, complement each other to improve recyclable tuning in this context.
Other approaches involve training additional domain-specific experts for new content domains. DEMix [91] addresses continual pre-training by incrementally training and integrating new experts (DEMix layer replacing every FFN layer in the transformer) for new domains. To ensure reasonable inference performance during testing when no domain information is available, DEMix proposes a parameter-free probabilistic approach, distinct from the gating function in MoE [259], to dynamically estimate a weighted mixture of domains. Introducing a new domain variable $D_{t}$ alongside each word $x_{t}$, the authors estimate the next word probability $p(x_{t}|{\bm{x}}_{<t})$ by marginalizing over all experts [^2] In the marginalization step of the following equation in the original paper, it is $p(D_{t}=j|{\bm{x}}_{t})$ instead of $p(D_{t}=j|{\bm{x}}_{<t})$, while we believe this is a minor typo, and hence here we use the updated version.:
$p(x_{t}|{\bm{x}}_{<t})$ $=\sum_{j=1}^{n}p(x_{t}|{\bm{x}}_{<t},D_{t}=j)\cdot p(D_{t}=j|{\bm{x}}_{<t})$
$=\sum_{j=1}^{n}p(x_{t}|{\bm{x}}_{<t},D_{t}=j)\cdot\left[\tfrac{p({\bm{x}}_{<t}|D_{t}=j)\cdot p(D_{t}=j)}{\sum_{j^{\prime}=1}^{n}p({\bm{x}}_{<t}|D_{t}=j^{\prime})\cdot p(D_{t}=j^{\prime})}\right],$ (12)
where all the probability terms $p(\cdot|\cdot,D_{t})$ conditioned on the domain variable $D_{t}$ are calculated by using a specific domain expert. The authors develop a large-scale continual pre-training evaluation benchmark comprising eight semantic domains for sequential training and another set of eight domains for assessing LLMs’ generalization ability. The DEMix framework’s modularization has been shown to facilitate efficient domain-adaptive pre-training, promote relevant knowledge during inference, and allow for removable components. Lifelong-MoE [46] shares a similar approach to DEMix [91] by incrementally training domain experts for each new domain. However, Lifelong-MoE differs in utilizing a token-level gating function to activate multiple experts for intermediate embedding calculation. During training, previous experts’ parameters and gating functions remain frozen, while knowledge distillation loss is employed to regulate parameter updates. Although the data distributions for evaluation are extremely large-scale (3 domains, 686 billion tokens in total), the Lifelong-MoE is able to efficiently mitigate the issue of horizontal forgetting.
It is noteworthy that some papers draw almost opposite conclusions regarding the significance of CPT. For instance, [55] continually pre-trains BERT [67] and RoBERTa [171] on five scientific domains and evaluates performance on downstream sentiment analysis. They observe that even baseline sequential pre-training does not exhibit severe forgetting, prompting reasonable questions about the necessity of continual pre-training or its suitable application scenarios.
Temporal Shift. In the context of continual learning amid content shifts, Multi-Task Learning (MTL) is often regarded as the upper limit achievable in continual learning scenarios [223, 288, 260]. However, this belief does not fully hold when considering continual learning under temporal shifts [120, 119, 68], as temporal shifts can introduce conflicting information, posing challenges for LLMs. For instance, the statement “Lionel Messi plays for team Barcelona” remains accurate from 2004 to 2021 but becomes false by 2024, as “Lionel Messi plays for team Inter Miami” becomes the correct statement.
Hence, as advocated by CKL [120] and TemporalWiki [119], LLMs undergoing continual adaptation to temporal shifts in the corpus must simultaneously achieve three objectives: (i) retention of old knowledge, (ii) acquisition of new knowledge, and (iii) update of the outdated knowledge (as a conflict resolution). They evaluate the same set of continual learning baseline methods, including parameter regularization (RecAdam [44]), rehearsal (Mix-review [102]), and parameter expansion (LoRA [110] and K-Adapter [290]), each highlighting distinct aspects of their impact. CKL [120] introduces a unified metric, FUAR (Forgetting / (Updated + Acquired) Ratio), to assess the three learning objectives collectively. They observe that parameter expansion consistently exhibits robust performance across all experimental conditions. In contrast, replay-based methods struggle to efficiently adapt to new knowledge acquisition and outdated knowledge update, leading to rapid forgetting of newly learned information during training. TemporalWiki [119], in contrast, constructs a series of temporal corpora and their differential sets from sequential snapshots of Wikipedia, investigating the efficacy of adapting LLMs to these differential sets. The study reveals that updating LLMs on these differential sets substantially enhances new knowledge acquisition and updates, requiring significantly less computational resources. Moreover, various continual learning techniques prove effective in mitigating horizontal forgetting during this process.
Additionally, LLPT [127] introduces temporal generalization evaluation for LLMs pre-trained on sequential corpora. Through experiments on a large-scale chronologically-ordered Tweet Stream, the authors demonstrate the superiority of CPT, combined with any continual learning technique, over a single task-specific LM, in terms of both knowledge acquisition and temporal generalization. Nonetheless, these preliminary experiments do not conclusively determine which specific continual learning method is more preferable than the others.
Another line of work, Temporal Language Models (TLMs), takes a different approach to address knowledge retention, acquisition, and update under temporal shifts by integrating temporal information into the learning process [244, 68, 271]. During training, they inject temporal information into training examples as prefixes of prompts, using special tokens [244], explicit year information [68], or syntax-guided structural information [271]. In sequential training experiments conducted by TempoT5 [68], comparison between continually and jointly pre-trained LMs demonstrates that CPT better balances adaptation and forgetting when the replay rate of past data is appropriately set.
The significance of addressing temporal shifts through continual pre-training is underscored by several industrial studies. For instance, [6] employs a dynamic vocabulary expansion algorithm and an efficient sub-sampling procedure to conduct CPT on large-scale emerging tweet data. Conversely, [175] adopts continual pre-training without explicit measures to constrain model updates, releasing a series of BERT-based LMs incrementally trained on new tweet data every three months. Preliminary experimental results demonstrate substantial improvements of continually pre-trained LMs over the base BERT model across downstream tasks. While some studies question the necessity of continually adapting LLMs along the temporal axis for environmental reasons, such as reducing CO2 emissions [9], the community commonly embraces CPT as a more efficient and environmentally friendly learning paradigm compared to the traditional “combine-and-retrain” approach.
Table 2: Summary of the existing studies that leverage Domain-Adaptive Pre-Training of LLMs, where the papers are organized in four main categories based on whether they (i) adopt the continual learning techniques and (ii) perform the evaluation for backward transfer (forgetting). In the column of Train Proc. (Training Process), we omit the phase of general Pre-Training. DAP represents Domain-Adaptive Pre-Training; SFT represents Supervised Fine-Tuning; IT represents Instruction Tuning. The prefix G- and D- represent General and Domain-Specific training process [166, 115], and the prefix U- represents them unified [310, 42]. The prefix MM- and LC- represents Multi-Modal and Long-Context training phases [185, 367, 245]. In the column of Continual Learning Eval., we consider two criteria: (i) Backward Transfer, i.e., performance degradation on the previous tasks, which is also known as catastrophic forgetting, (ii) Forward Transfer, i.e., the performance gained by DAP while transferring the LLMs to the downstream tasks. We use L and Perp. to denote Loss and Perplexity, FT to denote Fine-Tuning, ZS and FS to denote Zero-Shot and Few-Shot Accuracy, HE and LLM to denote the Human Evaluation and LLM Evaluation for generative tasks. Among 33 papers presented in this table that adopt DAP during the development, nearly 65% (22/33) of them explicitly study the influence of DAP from a continual learning perspective: they either evaluate the degree of forgetting, or adopt the continual learning techniques to prevent forgetting of the general knowledge. However, there is a significant lack of diversity of the continual learning techniques adopted in these works (only Replay and LoRA), which advocates the further study on the efficacy of vertical continual learning in the realm of LLMs.
Rehearsal Param. Reg. Arch. Exp. Backward Transfer Forward Transfer
Medical BioMedGPT [185] DAP $\rightarrow$ MM-SFT Llama2 FT
Financial BBT-Fin [177] DAP T5 FT
Financial CFGPT [152] DAP $\rightarrow$ SFT InternLM Q-LoRA${}_{\text{(SFT)}}$ HE1
Scientific AstroLlama [209] DAP LlaVa Perp.
Scientific OceanGPT [20] DAP $\rightarrow$ IT Vicuna Llama2-chat ChatGLM2 LoRA${}_{\text{(IT)}}$ HE
Scientific K2 [63] DAP $\rightarrow$ SFT Llama LoRA${}_{\text{(SFT)}}$ Perp. | ZS | LLM
Scientific MarineGPT [367] MM-DAP $\rightarrow$ MM-IT Llama HE
Code CodeGen [215] DAP $\rightarrow$ DAP CodeGen Perp. | ZS
Code Comment-Aug [269] IT $\rightarrow$ DAP Llama2 Code Llama InternLM2 ZS
EventTemporal EcoNet [93]1 DAP $\rightarrow$ FT BERT RoBERTa FT
CommonSense CALM [369] DAP $\rightarrow$ FT T5 FT
Medical [88] DAP $\rightarrow$ FT Llama2 FS | FT FS | FT
Financial [323] DAP Pythia L | FS L | FS
Scientific GeoGalactica [166] DAP $\rightarrow$ G-SFT $\rightarrow$ D-SFT GAL ZS Perp. | ZS | LLM
Code StarCoder [158] DAP StarCoder Perp. | ZS | FS Perp. | ZS | FS
Code DeepSeek-Coder [87] DAP DS-LLM ZS | FS ZS
Multi-Domain DAPT [92] DAP $\rightarrow$ FT RoBERTa Loss L | FT
Financial WeaverBird [326] DAP GLM2 LoRA HE
Code IRCoder [221] DAP StarCoder DS-Coder Code Llama LoRA ZS
Code Code Llama [245] DAP $\rightarrow$ LC-FT $\rightarrow$ IT DAP $\rightarrow$ DAP $\rightarrow$ LC-FT Llama2 Replay Perp. | ZS
Legal SaulLM [52] DAP $\rightarrow$ U-IT Mistral Replay Perp. | ZS
Medical PMC-Llama [311] DAP $\rightarrow$ IT Llama Replay ZS | FT
Scientific Llema [10] DAP Code Llama Replay Perp. | FS

### v2

@@ 4.1.3. Distributional Shifts in CPT
Content Shift. Without using complex CL techniques, (yildiz2024investigating, ) explores the large-scale CPT over 159 content domains, and makes key observations about the domain structures and model properties. It shows that CPT on various domains can effectively improve models’ adaptation ability compared to DAP on single domain. Similarly, (gupta2023continual, ) continues the pre-training phase of Pythia (biderman2023pythia, ) with no complex CL techniques. The study focuses on improving CPT with simple learning rate (re)warm-up. They discover that re-warming consistently improves models trained from scratch. Built upon this simple observation, (ibrahim2024simple, ) further shows that proper combination of learning rate re-warming and re-decay, and replay of the previous data is sufficient to achieve a comparable performance as full re-training.
Another pioneering work, LLPT (jin2022lifelong, ), establishes a comprehensive training and evaluation protocol for a series of content-level distributional shifts, referred to as “domain-incremental data streams”. They assess multiple CL methods and, similar to findings in (gogoulou2024continual, ), find consistent forward knowledge transfer, yet horizontal forgetting remains significant. Besides, contrary to the common understanding that experience replay (ER, (chaudhry2019tiny, )) is the most efficient approach to preventing forgetting, the authors find it ineffective in the case of CPT. They speculate that ER’s inefficiency may stem from overfitting issues (yuan2020revisiting, ; jin2022lifelong, ). Recyclable Tuning (qin2023recyclable, ) is the first study to consider both upstream LLM suppliers and downstream consumers at the same time. It shows that if the upstream supplier continually pre-trains LLMs, with or without replay, consumer-side efficiency can be boosted by recycling previously learned update components. Two CL techniques, initialization from outdated components and knowledge distillation, complement each other in improving adaptation efficiency in this context.
Other approaches involve training additional domain-specific experts for new content domains. DEMix (gururangan2022demix, ) addresses CPT by incrementally training and integrating new experts (DEMix layer) for new domains. To ensure reasonable inference performance during testing when no domain information is available, DEMix proposes a parameter-free probabilistic approach to dynamically estimate a weighted mixture of domains. Introducing a new domain variable $D_{t}$ alongside each word $x_{t}$, the authors estimate the next word probability $p(x_{t}|{\bm{x}}_{<t})$ by marginalizing over all experts [^2] In the marginalization step of the following equation in the original paper, it is $p(D_{t}=j|{\bm{x}}_{t})$ instead of $p(D_{t}=j|{\bm{x}}_{<t})$, while we believe this is a minor typo, and hence here we use the updated version.:
$p(x_{t}|{\bm{x}}_{<t})$ $=\sum_{j=1}^{n}p(x_{t}|{\bm{x}}_{<t},D_{t}=j)\cdot p(D_{t}=j|{\bm{x}}_{<t})=\sum_{j=1}^{n}p(x_{t}|{\bm{x}}_{<t},D_{t}=j)\cdot\left[\tfrac{p({\bm{x}}_{<t}|D_{t}=j)\cdot p(D_{t}=j)}{\sum_{j^{\prime}=1}^{n}p({\bm{x}}_{<t}|D_{t}=j^{\prime})\cdot p(D_{t}=j^{\prime})}\right],$
where the conditional probability terms $p(\cdot|\cdot,D_{t})$ are calculated by using a specific domain expert. The DEMix framework’s modularization has been shown to facilitate efficient domain-adaptive pre-training, promote relevant knowledge during inference, and allow for removable components. Lifelong-MoE (chen2023lifelong, ), similar to DEMix (gururangan2022demix, ), incrementally trains domain experts for new domains. However, Lifelong-MoE differs from DEMix in utilizing a token-level gating function to activate multiple experts for intermediate embedding calculation. During training, previous experts’ parameters and gating functions remain frozen, and knowledge distillation loss is employed to regulate parameter updates, which thereby makes Lifelong-MoE robust against the issue of horizontal forgetting.
It is noteworthy that some papers draw almost opposite conclusions regarding the significance of CPT for content shifts. For instance, (cossu2022continual, ) continually pre-trains BERT-based models (devlin2018bert, ; liu2019roberta, ) on five scientific domains and evaluates performance on downstream sentiment analysis. They observe that even the trivial sequential pre-training does not exhibit severe forgetting, prompting reasonable questions about the necessity of CPT.
Temporal Shift. In the context of CPT amid content shifts, Multi-Task Learning (MTL) is often regarded as the upper bound achievable (pentina2016theoretical, ; wang2024comprehensive, ; shi2024unified, ). However, this belief does not fully hold when considering CL under temporal shifts (jang2022towards, ; jang2022temporalwiki, ; dhingra2022time, ), as temporal shifts can introduce conflicting information, posing challenges for LLMs. For instance, the statement “Lionel Messi plays for team Barcelona” remains accurate from 2004 to 2021 but becomes false by 2024, as “Lionel Messi plays for team Inter Miami” becomes the correct statement.
Hence, as advocated by CKL (jang2022towards, ) and TemporalWiki (jang2022temporalwiki, ), LLMs undergoing continual adaptation to temporal shifts must simultaneously achieve three objectives: (i) retention of old knowledge, (ii) acquisition of new knowledge, and (iii) update of the outdated knowledge. They evaluate the same set of continual learning baseline methods (chen2020recall, ; he2021analyzing, ; hu2022lora, ; wang2021kadapter, ), each highlighting distinct aspects of their impact. CKL (jang2022towards, ) observes that parameter expansion consistently exhibits robust performance across all experimental conditions. In contrast, replay-based methods struggle to efficiently adapt to new knowledge acquisition and outdated knowledge update, leading to rapid forgetting of newly learned information during training. TemporalWiki (jang2022temporalwiki, ) constructs a series of temporal corpora and their differential sets from sequential snapshots of Wikipedia, revealing that updating LLMs on these differential sets substantially enhances new knowledge acquisition and updates, requiring significantly less computational resources, and various CL techniques prove effective in mitigating horizontal forgetting during this process. LLPT (jin2022lifelong, ) introduces temporal generalization evaluation for LLMs pre-trained on sequential corpora. Through experiments on a large-scale chronologically-ordered Tweet Stream, the authors demonstrate the superiority of CPT combined with CL techniques to task-specific LMs, in terms of both knowledge acquisition and temporal generalization. Nonetheless, these preliminary experiments do not conclusively determine which specific CL method is more preferable than the others.
Another line of work, Temporal Language Models (TLMs), takes a different approach to address knowledge retention, acquisition, and update under temporal shifts by integrating temporal information into the model (rosin2022time, ; dhingra2022time, ; su2023efficient, ). During training, they inject temporal information into training examples as prefixes of prompts, using special tokens (rosin2022time, ), explicit year information (dhingra2022time, ), or syntax-guided structural information (su2023efficient, ). In sequential training experiments conducted by TempoT5 (dhingra2022time, ), comparison between continually and jointly pre-trained LMs demonstrates that CPT better balances adaptation and forgetting when the replay rate of past data is appropriately set.
Others. CPT as a technique to progressively attain novel knowledge, can be used to refine LLMs’ behavior. CEM (zhao2024large, ) collects examples where the model’s response is incorrect and continually trains the model on these, along with a supplemental dataset. RHO-1 (lin2024rho, ) proposes Selective Language Modeling (SLM), which employs a reference model to evaluate the perplexity of each token in the training corpus, and continually pre-trains the model on high-perplexity tokens. Similarly, IR-DRO (chen2024take, ) re-trains the model on re-weighted examples from the original pre-training dataset, focusing more on higher-loss sequences.
The significance of addressing temporal shifts through CPT is underscored by several industrial studies. For instance, (amba2021dynamic, ) employs a dynamic vocabulary expansion algorithm and an efficient sub-sampling procedure to conduct CPT on large-scale emerging tweet data. Conversely, (loureiro2022timelms, ) adopts CPT without explicit measures to constrain model updates, releasing a series of BERT-based LMs incrementally trained on new tweet data every three months. Preliminary experimental results demonstrate substantial improvements of continually pre-trained LMs over the base BERT model across downstream tasks. While some studies question the necessity of continually adapting LLMs along the temporal axis for environmental reasons, such as reducing CO2 emissions (attanasio2023worth, ), the community commonly embraces CPT as a more efficient learning paradigm compared to the traditional “combine-and-retrain” approach.
4.2. Domain-Adaptive Pre-training (DAP)
Background of DAP. Institutions, regardless of size, often possess significant amounts of unlabeled, domain-specific data. This data bridges the gap between general-purpose LLMs trained on diverse corpora and fine-tuned LLMs designed for specific downstream tasks. Leveraging this data as a preparatory stage can facilitate effective adaptation of LLMs to downstream tasks. Such process of “continued/continual/continuous pre-training” (yan2023af, ; guo2023continuous, ; ma2023ecomgptct, ; han2021econet, ; xie2023efficient, ; xie2023quert, ; huang2023lawyer, ; Lu2023BBTFin, ; Xie2023PIXIU, ; Azerbayev2023LLEMMA, ; yue2023mammoth, ; colombo2024saullm7b, ; Zhang2024SciGLM, ; shen2024tag, ), “further pre-training” (song2024code, ; lin2023geogalactica, ; deng2023learning, ; Rubungo2023LLM-Prop, ; agarwal2024structured, ), “domain tuning” (rongali2021continual, ), “knowledge enhancement pre-training” (Lu2023BBTFin, ), and “knowledge injection training” (wu2023pmc, ) is unified and termed “Domain Adaptive Pre-training (DAP)” (gururangan2020dont, ) for clarity and consistency throughout this survey. In the pioneering work of domain-adaptive pre-training (DAPT) (gururangan2020dont, ), the authors continuously pre-train the language models on a larger domain-specific dataset before fine-tuning them to the downstream tasks, resulting in universally improved performance aross various tasks. As the observation above has been validated on multiple domains in parallel, including BioMed, CS, News, and Reviews (gururangan2020dont, ), practitioners commonly accept that employing DAP on additional unlabeled domain-specific data benefits downstream tasks. Consequently, this technique has become widely deployed in many modern LLMs.
Summary of LLMs with DAP. We provide a summary of existing studies utilizing DAP for LLMs in Table 2. Each entry is characterized by three main features: (i) training process specifications, encompassing the vertical domain for which LLMs are trained, the training pipeline preceding release, and the LLM architecture employed; (ii) adopted continual learning techniques, including rehearsal, parameter regularization, and architecture expansion; and (iii) evaluation metrics for CL, such as backward transfer (forgetting) and forward transfer (adaptation to downstream data).
4.2.1. General Observation on DAP
Several key observations emerge regarding the research landscape of DAP (Table 2).
4.2.2. Different Domains of DAP
Legal Domain. Given the legal industry’s demand for managing ever-growing volumes of legal documents, there’s a burgeoning need to harness LLMs to aid legal professionals in navigating, interpreting, and generating high-quality legal materials (xiao2021lawformer, ; savelka2023explaining, ; yue2023disc, ). In Layer Llama (huang2023lawyer, ), the authors gathered publicly available legal texts from China Courts websites, totaling approximately 10 billion tokens as noted in a GitHub issue (lawyerllam-git, ). In SaulLM (colombo2024saullm7b, ), the authors collected the DAP corpus from various jurisdictions in different countries, resulting in a corpus of 30 billion tokens to cover diverse aspects of legal texts. When combined with previously available datasets (gao2020pile, ; koehn2005europarl, ), the total number of tokens used for legal-domain DAP reaches 94 billion. The substantial volume of DAP data, while offering valuable insights into specific domains, increases the risk of vertical forgetting of the general knowledge due to the large number of update steps involved. To mitigate this issue, SaulLM incorporates general data from Wikipedia, StackExchange, and GitHub into the DAP data, constituting about 2% of the final dataset (colombo2024saullm7b, ). Similarly, Lawyer Llama incorporates replaying general-domain data during DAP, but the replay rate is not disclosed (huang2023lawyer, ). (takahashi2024pretraining, ) also replays of non-latest business documents during DAP when building a Japanese business-specific LLM.
Medical Domain. The development of LLMs holds promise for revolutionary changes in the medical industry, offering potential improvements in efficiency and quality across medical communication, disease diagnosis, and decision-making for doctors (he2024foundation, ; li2023chatdoctor, ; singhal2023large, ; jeblick2022chatgpt, ; chen2023utility, ). Efforts have been made to develop medical specialists by either training an LLM from scratch (Han2023MedAlpaca, ; Singhal2023MedPaLM2, ; gu2021domain, ; luo2022biogpt, ) or fine-tuning publicly-available LLMs to meet specific medical needs (luo2023biomedgpt, ; wu2023pmc, ; Chen2023HuatuoGPTII, ; xiong2023doctorglm, ; bao2023discmedllm, ; zhang2023alpacare, ). Among these approaches, DAP techniques have been extensively utilized to preserve the communication and instruction-following abilities of a general LLM, preparing it for subsequent medical applications (luo2023biomedgpt, ; wu2023pmc, ; Chen2023HuatuoGPTII, ).
BioMedGPT (luo2023biomedgpt, ) is a multi-modal biomedical language model that integrates representations of human language and the language of life (molecules, proteins, cells, genes, etc.). Prior to final multi-modal supervised fine-tuning, the authors initialize the model from Llama2-Chat (touvron2023llama2, ) and conduct DAP using extensive biomedical documents from S2ORC (lo2020s2orc, ), without considering any CL techniques or evaluations. In (guo2023continuous, ), DAP is performed using Chinese medical encyclopedias and online expert articles, with next-token prediction as the training objective. During DAP, the performance gradually deteriorates on general-domain datasets as the training step increases, but improves on the downstream medical examination tasks (hendryckstest2021, ; liu2023benchmarking, ; li2023cmmlu, ). PMC-LLama (wu2023pmc, ) gathers biomedical papers from S2ORC (lo2020s2orc, ) and medical textbooks for “knowledge injection training”. During this phase, a general language corpus from RedPajama-Data (together2023redpajama, ) is replayed at a 5% rate within a training batch. However, the paper does not analyze the effectiveness of this operation of mixing in general-domain data for DAP.
To mitigate vertical forgetting, AF Adapter (yan2023af, ) proposes an adapter structure extending the width of Attention layers and FFNs for acquiring domain knowledge and only the adapters are tuned during DAP. Similarly, Hippocrates (acikgoz2024hippocrates, ) deploys LoRA during DAP to both have medical-specific knowledge injected and general ability preserved. Me-Llama (xie2024me, ) mixes in about 25% of the general-domain data for DAP on the clinical notes and biomedical articles, which achieves even positive backward transfer on MMLU (hendryckstest2021, ). HuatuoGPT-II (Chen2023HuatuoGPTII, ) proposes to fuse the DAP into the final SFT, making the two-stage development one unified process. The challenge of such process mainly comes from the data heterogeneity of DAP’s unlabeled corpus. The authors address this challenge by reformulating paragraphs of data into (instruction, output) format using existing large language models. They further employ a priority sampling strategy to avoid compromising downstream ability, a pitfall observed in the fixed-rate data mixing strategy (touvron2023llama2, ). This paper empirically demonstrates the superiority of unified one-stage SFT over two-stage training, questioning the reasonability of the current DAP. On medical-domain data, (rongali2021continual, ) finds that LMs constrained by CL techniques on source domains exhibit greater robustness to future domain shifts. Specifically, they identify that parameter regularization techniques like EWC (kirkpatrick2017overcoming, ), despite slightly higher cost, can facilitate positive forward and backward transfer.
Table 2. Summary of the existing studies that leverage Domain-Adaptive Pre-training of LLMs, where the papers are organized in four main categories based on whether they (i) adopt the continual learning techniques and (ii) perform the evaluation for backward transfer (forgetting). In the column of Train Proc. (Training Process), we omit the phase of general Pre-Training. DAP represents Domain-Adaptive Pre-Training; SFT represents Supervised Fine-Tuning; IT represents Instruction Tuning. The prefix G- and D- represent General and Domain-Specific training process (lin2023geogalactica, ; huang2023lawyer, ), and the prefix U- represents them unified (wu2024llama, ; Chen2023HuatuoGPTII, ). The prefix MM- and LC- represents Multi-Modal and Long-Context training phases (luo2023biomedgpt, ; Zheng2023MarineGPT, ; rozière2024code, ). In the column of Continual Learning Eval., we consider two criteria: (i) Backward Transfer, i.e., performance degradation on the previous tasks, which is also known as catastrophic forgetting, (ii) Forward Transfer, i.e., the performance gained by DAP while transferring the LLMs to the downstream tasks. We use L and Perp. to denote Loss and Perplexity, FT to denote Fine-Tuning, ZS and FS to denote Zero-Shot and Few-Shot Accuracy, HE and LLM to denote the Human Evaluation and LLM Evaluation for generative tasks.
Rehearsal Param. Reg. Arch. Exp. Backward Transfer Forward Transfer
Medical BioMedGPT (luo2023biomedgpt, ) DAP $\rightarrow$ MM-SFT Llama2 FT
Financial BBT-Fin (Lu2023BBTFin, ) DAP T5 FT
Financial CFGPT (li2023cfgpt, ) DAP $\rightarrow$ SFT InternLM Q-LoRA${}_{\text{(SFT)}}$ HE1
Scientific AstroLlama (Nguyen2023AstroLLaMA, ) DAP LlaVa Perp.
Scientific OceanGPT (Bi2023OCEANGPT, ) DAP $\rightarrow$ IT Vicuna Llama2-chat ChatGLM2 LoRA${}_{\text{(IT)}}$ HE
Scientific K2 (deng2023learning, ) DAP $\rightarrow$ SFT Llama LoRA${}_{\text{(SFT)}}$ Perp. | ZS | LLM
Scientific MarineGPT (Zheng2023MarineGPT, ) MM-DAP $\rightarrow$ MM-IT Llama HE
Code CodeGen (nijkamp2022codegen, ) DAP $\rightarrow$ DAP CodeGen Perp. | ZS
Code Comment-Aug (song2024code, ) IT $\rightarrow$ DAP Llama2 Code Llama InternLM2 ZS
EventTemporal EcoNet (han2021econet, )1 DAP $\rightarrow$ FT BERT RoBERTa FT
CommonSense CALM (zhou2020pre, ) DAP $\rightarrow$ FT T5 FT
Multi-Domain BLADE (li2024blade, ) DAP $\rightarrow$ IT BLOOMZ ZS
Scientific ClimateGPT (thulke2024climategpt, ) DAP $\rightarrow$ IT $\rightarrow$ RAG Llama2 FS | Ret.
Medical (guo2023continuous, ) DAP $\rightarrow$ FT Llama2 FS | FT FS | FT
Financial (xie2023efficient, ) DAP Pythia L | FS L | FS
Scientific GeoGalactica (lin2023geogalactica, ) DAP $\rightarrow$ G-SFT $\rightarrow$ D-SFT GAL ZS Perp. | ZS | LLM
Code StarCoder (li2023starcoder, ) DAP StarCoder Perp. | ZS | FS Perp. | ZS | FS
Code DeepSeek-Coder (guo2024deepseekcoder, ) DAP DeepSeek-LLM ZS | FS ZS
Multi-Domain DAPT (gururangan2020dont, ) DAP $\rightarrow$ FT RoBERTa Loss L | FT
Financial WeaverBird (Xue2023WeaverBird, ) DAP GLM2 LoRA HE
Code IRCoder (paul2024ircoder, ) DAP StarCoder DeepSeek-Coder Code Llama LoRA ZS
Code Code Llama (rozière2024code, ) DAP $\rightarrow$ LC-FT $\rightarrow$ IT DAP $\rightarrow$ DAP $\rightarrow$ LC-FT Llama2 Replay Perp. | ZS
Legal SaulLM (colombo2024saullm7b, ) DAP $\rightarrow$ U-IT Mistral Replay Perp. | ZS
Medical PMC-Llama (wu2023pmc, ) DAP $\rightarrow$ IT Llama Replay ZS | FT
Scientific Llema (Azerbayev2023LLEMMA, ) DAP Code Llama Replay Perp. | FS